OpenStudy (anonymous):

Find the equation of the circle whose center and radius are given. center ( 5, 6), radius = 3

OpenStudy (anonymous):

The general form of the circle equation is (x - p)^2 + (y - q)^2 = r^2 where (p, q) is the center and r is the radius

OpenStudy (anonymous):

(x - 5)^2 + (y - 6)^2 = 9